# Rethinking Probabilistic Circuit Parameter Learning

### v1

@@ Acknowledgements.
This work was funded in part by the DARPA ANSR, CODORD, and SAFRON programs under awards FA8750-23-2-0004, HR00112590089, and HR00112530141, NSF grant IIS1943641, and gifts from Adobe Research, Cisco Research, and Amazon. Approved for public release; distribution is unlimited.
@@ Appendix A Structural Properties of PCs
We provide formal definitions of smoothness and decomposability. Please refer to Choi et al. [2020] for a comprehensive overview.
@@
Define the scope $\mathtt{Var}(n)$ of a PC node $n$ as the set of all variables defined by its descendant input nodes. A PC ${p}$ is smooth if for every sum node $n$, its children share the same scope: $\forall c_{1},c_{2}\in\mathsf{ch}(n)$, $\mathtt{Var}(c_{1})=\mathtt{Var}(c_{2})$. ${p}$ is decomposable if for every product node $n$, its children have disjoint scopes: $\forall c_{1},c_{2}\in\mathsf{ch}(n)$ ($c_{1}\neq c_{2}$), $\mathtt{Var}(c_{1})\cap\mathtt{Var}(c_{2})=\varnothing$.
@@ B.1 Interpreting the EM Algorithm of PCs
${p}_{n}(\boldsymbol{x},\mathbf{Z}_{n}=i,\boldsymbol{z}_{\backslash n})=\sum_{c\in\mathsf{ch}(n)}\exp(\phi_{n,c})\cdot{p}_{c}(\boldsymbol{x},\mathbf{Z}_{n}=i,\boldsymbol{z}_{\backslash n})\cdot\mathbbm{1}[c=c_{i}],$
where we define $c_{i}$ as the $i$-th child node of $n$, and $\mathbf{Z}_{\backslash n}:=\mathbf{Z}\backslash Z_{n}$.
@@
$\mathtt{supp}(n):=\begin{cases}\{(\boldsymbol{x},\boldsymbol{z}):{p}_{n}(\boldsymbol{x})>0\}&n\text{\leavevmode\nobreak\ is\leavevmode\nobreak\ an\leavevmode\nobreak\ input\leavevmode\nobreak\ node},\\ \bigcap_{c\in\mathsf{ch}(n)}\mathtt{supp}(c)&n\text{\leavevmode\nobreak\ is\leavevmode\nobreak\ a\leavevmode\nobreak\ product\leavevmode\nobreak\ node},\\ \bigcup_{c\in\mathsf{ch}(n)}\left(\{(\boldsymbol{x},\boldsymbol{z}):z_{n}=c\}\cap\mathsf{supp}(c)\right)&n\text{\leavevmode\nobreak\ is\leavevmode\nobreak\ a\leavevmode\nobreak\ sum\leavevmode\nobreak\ node},\end{cases}$
@@
$S_{\phi_{n,c}}(\boldsymbol{x},\boldsymbol{z})=\mathbbm{1}\Big{[}(\boldsymbol{x},\boldsymbol{z})\in\mathtt{supp}(c)\text{\leavevmode\nobreak\ and\leavevmode\nobreak\ }z_{n}=c\Big{]},$
@@
The following part partially follows Kunstner et al. [2021]. We plug in the exponential family distribution form of the PC into the definition of $Q_{\boldsymbol{\phi}}(\boldsymbol{\phi}^{\prime})$:
@@
$=\frac{1}{|\mathcal{D}|}\sum_{\boldsymbol{x}\in\mathcal{D}}\sum_{\boldsymbol{z}}{p}_{\boldsymbol{\phi}}(\boldsymbol{z}|\boldsymbol{x})\left[\left\langle S(\boldsymbol{x},\boldsymbol{z}),\boldsymbol{\phi}^{\prime}\right\rangle-A(\boldsymbol{\phi}^{\prime})\right],$ Definition of $\triangleright\text{Definition\leavevmode\nobreak\ of\leavevmode\nobreak\ }{p}_{\boldsymbol{\phi}^{\prime}}(\boldsymbol{x},\boldsymbol{z})$
[…]
@@
$=\left\langle\frac{\partial A(\boldsymbol{\phi})}{\partial\boldsymbol{\phi}},\boldsymbol{\phi}-\boldsymbol{\phi}^{\prime}\right\rangle+A(\boldsymbol{\phi}^{\prime})-A(\boldsymbol{\phi}),$ Since $\triangleright\text{Since\leavevmode\nobreak\ }\frac{\partial A(\boldsymbol{\phi})}{\partial\boldsymbol{\phi}}=\mathbb{E}_{\boldsymbol{x},\boldsymbol{z}\sim{p}_{\boldsymbol{\phi}}}\left[S(\boldsymbol{x},\boldsymbol{z})\right]$
@@
$\frac{1}{|\mathcal{D}|}\sum_{\boldsymbol{x}\in\mathcal{D}}\log{p}_{\boldsymbol{\phi}}(\boldsymbol{x})+\left\langle\frac{\partial\log{p}_{\boldsymbol{\phi}}(\boldsymbol{x})}{\partial\boldsymbol{\phi}},\boldsymbol{\phi}^{\prime}-\boldsymbol{\phi}\right\rangle-\mathtt{KL}_{\boldsymbol{\phi}}(\boldsymbol{\phi}^{\prime}).$ (9)
@@
$Z_{n}(\boldsymbol{\phi})=\begin{cases}1&n\text{\leavevmode\nobreak\ is\leavevmode\nobreak\ an\leavevmode\nobreak\ input\leavevmode\nobreak\ node},\\ \prod_{c\in\mathsf{ch}(n)}Z_{c}(\boldsymbol{\phi})&n\text{\leavevmode\nobreak\ is\leavevmode\nobreak\ a\leavevmode\nobreak\ product\leavevmode\nobreak\ node},\\ \sum_{c\in\mathsf{ch}(n)}\exp(\phi_{n,c})\cdot Z_{c}(\boldsymbol{\phi})&n\text{\leavevmode\nobreak\ is\leavevmode\nobreak\ a\leavevmode\nobreak\ sum\leavevmode\nobreak\ node}.\end{cases}$
@@
$=\sum_{c^{\prime}\in\mathsf{ch}(n)}\exp(\phi_{n,c^{\prime}}),$ Since we assume $\triangleright\text{Since\leavevmode\nobreak\ we\leavevmode\nobreak\ assume\leavevmode\nobreak\ }\forall c,Z_{c}(\boldsymbol{\phi})=1$
@@
$=\mathtt{TD}_{m}(n)\cdot\exp(\phi_{n,c})+\sum_{c^{\prime}\in\mathsf{ch}(n),c^{\prime}\neq c}\exp(\phi_{n,c^{\prime}}),$ Since $\triangleright\text{Since\leavevmode\nobreak\ }\forall c,\mathtt{TD}_{c}(c)=1$
@@
$=\sum_{b\in\mathcal{A}}\exp(\phi_{m,b})\cdot\mathtt{TD}_{b}(n).$ Since $\triangleright\text{Since\leavevmode\nobreak\ }\mathtt{TD}_{m}(b)=\exp(\phi_{m,b})$
@@
We move on to the second equality in Lemma 1. According to Vergari et al. [2021], $\mathtt{KL}_{\boldsymbol{\phi}}(\boldsymbol{\phi}^{\prime})$ can be computed recursively as follows (define $\mathtt{KL}_{\boldsymbol{\phi}}^{n}(\boldsymbol{\phi}^{\prime})$ as the KLD w.r.t. $n$):
$\mathtt{KL}_{\boldsymbol{\phi}}^{n}(\boldsymbol{\phi}^{\prime})=\begin{cases}0&n\text{\leavevmode\nobreak\ is\leavevmode\nobreak\ an\leavevmode\nobreak\ input\leavevmode\nobreak\ node},\\ \sum_{c\in\mathsf{ch}(n)}\mathtt{KL}_{\boldsymbol{\phi}}^{c}(\boldsymbol{\phi}^{\prime})&n\text{\leavevmode\nobreak\ is\leavevmode\nobreak\ a\leavevmode\nobreak\ product\leavevmode\nobreak\ node},\\ \sum_{c\in\mathsf{ch}(n)}\exp(\phi_{n,c})\big{(}\phi_{n,c}-\phi^{\prime}_{n,c}\big{)}+\exp(\phi_{n,c})\cdot\mathtt{KL}_{\boldsymbol{\phi}}^{c}(\boldsymbol{\phi}^{\prime})&n\text{\leavevmode\nobreak\ is\leavevmode\nobreak\ a\leavevmode\nobreak\ sum\leavevmode\nobreak\ node}.\end{cases}$ (14)
@@
$\mathtt{KL}_{\boldsymbol{\phi}}^{n}(\boldsymbol{\phi}^{\prime})=\begin{cases}0&n\text{\leavevmode\nobreak\ is\leavevmode\nobreak\ an\leavevmode\nobreak\ input\leavevmode\nobreak\ node},\\ \sum_{c\in\mathsf{ch}(n)}\mathtt{KL}_{\boldsymbol{\phi}}^{c}(\boldsymbol{\phi}^{\prime})&n\text{\leavevmode\nobreak\ is\leavevmode\nobreak\ a\leavevmode\nobreak\ product\leavevmode\nobreak\ node},\\ \sum_{c\in\mathsf{ch}(n)}\exp(\phi_{n,c})\cdot\mathtt{KL}_{\boldsymbol{\phi}}^{c}(\boldsymbol{\phi}^{\prime})&n\text{\leavevmode\nobreak\ is\leavevmode\nobreak\ a\leavevmode\nobreak\ sum\leavevmode\nobreak\ node}.\end{cases}$
@@ B.3 Decomposition of Parameter Flows
$=\frac{\theta_{n,c}}{\tilde{{p}}_{\boldsymbol{\phi}}^{n}(\boldsymbol{x})}\cdot\frac{\partial\tilde{{p}}_{\boldsymbol{\phi}}^{n}(\boldsymbol{x})}{\partial\theta_{n,c}},$ By definition $\triangleright\text{By\leavevmode\nobreak\ definition\leavevmode\nobreak\ }\theta_{n,c}=\exp(\phi_{n,c})$
@@ Appendix C Global Parameter Renormalization of PCs
$\forall\boldsymbol{x},\;{p}_{\boldsymbol{\phi}^{\prime}}^{n}(\boldsymbol{x})=\frac{1}{Z_{n}(\boldsymbol{\phi})}\cdot{p}_{\boldsymbol{\phi}}^{n}(\boldsymbol{x}),$
where $Z(\boldsymbol{\phi}):=\sum_{\boldsymbol{x}}{p}_{\boldsymbol{\phi}}^{n}(\boldsymbol{x})$ is the partition function of ${p}_{\boldsymbol{\phi}}^{n}$.
The Algorithm.
First, we perform a feedforward evaluation of the PC to compute the partition function $Z_{n}(\boldsymbol{\phi})$ of every node $n$:
$Z_{n}(\boldsymbol{\phi})=\begin{cases}1&n\text{\leavevmode\nobreak\ is\leavevmode\nobreak\ an\leavevmode\nobreak\ input\leavevmode\nobreak\ node},\\ \prod_{c\in\mathsf{ch}(n)}Z_{c}(\boldsymbol{\phi})&n\text{\leavevmode\nobreak\ is\leavevmode\nobreak\ a\leavevmode\nobreak\ product\leavevmode\nobreak\ node},\\ \sum_{c\in\mathsf{ch}(n)}\exp(\phi_{n,c})\cdot Z_{c}(\boldsymbol{\phi})&n\text{\leavevmode\nobreak\ is\leavevmode\nobreak\ a\leavevmode\nobreak\ sum\leavevmode\nobreak\ node}.\end{cases}$
Next, for every sum edge $(n,c)$ (i.e., $n$ is a sum node and $c$ is one of its children), we update the parameter as
$\phi^{\prime}_{n,c}=\log\left(\frac{\theta_{n,c}\cdot Z_{\boldsymbol{\phi}}(c)}{Z_{\boldsymbol{\phi}}(n)}\right),$ (16)
where $\theta_{n,c}:=\exp(\phi_{n,c})$. The existence of this normalization algorithm has been previously shown by Martens and Medabalimi [2014], although they did not provide an easy-to-implement algorithm.
@@
We begin by proving the correctness of the algorithm. Specifically, we show by induction that ${p}_{\boldsymbol{\phi}^{\prime}}^{n}(\boldsymbol{x})={p}_{\boldsymbol{\phi}}^{n}(\boldsymbol{x})/Z_{\boldsymbol{\phi}}(n)$ for every $n$ and $\boldsymbol{x}$. In the base case, all input nodes satisfy the equation since they are assumed to be normalized. Next, given a product node $n$, assume the distributions encoded by all its children $c$ satisfy
$\forall c\in\mathsf{ch}(n),\;{p}_{\boldsymbol{\phi}^{\prime}}^{c}(\boldsymbol{x})={p}_{\boldsymbol{\phi}}^{c}(\boldsymbol{x})/Z_{\boldsymbol{\phi}}(c).$ (17)
@@
${p}_{\boldsymbol{\phi}^{\prime}}^{n}(\boldsymbol{x})$ $=\prod_{c\in\mathsf{ch}(n)}{p}_{\boldsymbol{\phi}^{\prime}}^{c}(\boldsymbol{x})=\prod_{c\in\mathsf{ch}(n)}{p}_{\boldsymbol{\phi}}^{c}(\boldsymbol{x})/Z_{\boldsymbol{\phi}}(c),$
$=\frac{\prod_{c\in\mathsf{ch}(n)}{p}_{\boldsymbol{\phi}}^{c}(\boldsymbol{x})}{\prod_{c\in\mathsf{ch}(n)}Z_{\boldsymbol{\phi}}(c)},$
$=\frac{{p}_{\boldsymbol{\phi}}^{n}(\boldsymbol{x})}{Z_{\boldsymbol{\phi}}(n)}.$
Finally, consider a sum node $n$ whose children satisfy Equation 17. We simplify ${p}_{\boldsymbol{\phi}^{\prime}}^{n}(\boldsymbol{x})$ in the following:
${p}_{\boldsymbol{\phi}^{\prime}}^{n}(\boldsymbol{x})$ $=\sum_{c\in\mathsf{ch}(n)}\theta^{\prime}_{n,c}\cdot{p}_{\boldsymbol{\phi}^{\prime}}^{c}(\boldsymbol{x}),$
@@
[…]
$=\frac{\sum_{c\in\mathsf{ch}(n)}\theta_{n,c}\cdot{p}_{\boldsymbol{\phi}}^{c}(\boldsymbol{x})}{Z_{\boldsymbol{\phi}}(n)},$
$={p}_{\boldsymbol{\phi}}^{n}(\boldsymbol{x})/Z_{\boldsymbol{\phi}}(n).$
@@
$\forall\boldsymbol{x},\;\frac{\partial\log{p}_{\boldsymbol{\phi}^{\prime}}(\boldsymbol{x})}{\partial\phi^{\prime}_{n,c}}=\frac{\partial\log{p}_{\boldsymbol{\phi}}(\boldsymbol{x})}{\partial\phi_{n,c}}.$
@@
$\forall\boldsymbol{x},\;\frac{\partial\log{p}_{\boldsymbol{\phi}^{\prime}}^{n}(\boldsymbol{x})}{\partial\log{p}_{\boldsymbol{\phi}^{\prime}}^{c}(\boldsymbol{x})}=\frac{\partial\log{p}_{\boldsymbol{\phi}}^{n}(\boldsymbol{x})}{\partial\log{p}_{\boldsymbol{\phi}}^{c}(\boldsymbol{x})}.$
@@
$\log{p}_{\boldsymbol{\phi}^{\prime}}^{n}(\boldsymbol{x})=\sum_{c\in\mathsf{ch}(n)}\log{p}_{\boldsymbol{\phi}^{\prime}}^{c}(\boldsymbol{x}).$
Hence, its derivative w.r.t. $\log{p}_{\boldsymbol{\phi}^{\prime}}^{c}(\boldsymbol{x})$ is $1$.

### v2

@@ Acknowledgements
This work was funded in part by the National University of Singapore under its Start-up Grant (Award No: SUG-251RES2505), the DARPA ANSR, CODORD, and SAFRON programs under awards FA8750-23-2-0004, HR00112590089, and HR00112530141, NSF grant IIS1943641, and gifts from Adobe Research, Cisco Research, Qualcomm, and Amazon. Approved for public release; distribution is unlimited.
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes]
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. [Yes]
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. [Yes/No/Not Applicable]
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. [Yes]
    2. (b)
      Complete proofs of all theoretical results. [Yes]
    3. (c)
      Clear explanations of any assumptions. [Yes]
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). [Yes]
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). [Yes]
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). [No]
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). [Yes]
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. [Yes]
    2. (b)
      The license information of the assets, if applicable. [Yes]
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. [Not Applicable]
    4. (d)
      Information about consent from data providers/curators. [Yes]
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. [Not Applicable]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. [Not Applicable]
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. [Not Applicable]
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. [Not Applicable]
Rethinking Probabilistic Circuit Parameter Learning:
Supplementary Materials
@@ Appendix A Structural Properties of PCs
We provide formal definitions of smoothness and decomposability. Please refer to Choi et al., (2020) for a comprehensive overview.
@@
Define the scope $\mathtt{Var}(n)$ of a PC node $n$ as the set of all variables defined by its descendant input nodes. A PC p{p} is smooth if for every sum node $n$, its children share the same scope: $\forall c_{1},c_{2}\in\mathsf{ch}(n)$, $\mathtt{Var}(c_{1})=\mathtt{Var}(c_{2})$. p{p} is decomposable if for every product node $n$, its children have disjoint scopes: $\forall c_{1},c_{2}\in\mathsf{ch}(n)$ ($c_{1}\neq c_{2}$), $\mathtt{Var}(c_{1})\cap\mathtt{Var}(c_{2})=\varnothing$.
@@ B.1 Interpreting the EM Algorithm of PCs
[…]
@@
[…]
@@
$S_{\phi_{n,c}}(\boldsymbol{x},\boldsymbol{z})=\mathbbm{1}\Big[(\boldsymbol{x},\boldsymbol{z})\in\mathtt{supp}(c)\text{~and~}z_{n}=c\Big],$
@@
The following part partially follows Kunstner et al., (2021). We plug in the exponential family distribution form of the PC into the definition of $Q_{\boldsymbol{\phi}}(\boldsymbol{\phi}^{\prime})$:
@@
$=\frac{1}{|\mathcal{D}|}\sum_{\boldsymbol{x}\in\mathcal{D}}\sum_{\boldsymbol{z}}{p}_{\boldsymbol{\phi}}(\boldsymbol{z}|\boldsymbol{x})\left[\left\langle S(\boldsymbol{x},\boldsymbol{z}),\boldsymbol{\phi}^{\prime}\right\rangle-A(\boldsymbol{\phi}^{\prime})\right],$ $\triangleright\text{Definition~of~}{p}_{\boldsymbol{\phi}^{\prime}}(\boldsymbol{x},\boldsymbol{z})$
$=\frac{1}{|\mathcal{D}|}\sum_{\boldsymbol{x}\in\mathcal{D}}\left\langle\mathbb{E}_{{p}_{\boldsymbol{\phi}}(\boldsymbol{z}|\boldsymbol{x})}\left[S(\boldsymbol{x},\boldsymbol{z})\right],\boldsymbol{\phi}^{\prime}\right\rangle-A(\boldsymbol{\phi}^{\prime}).$ $\triangleright\text{Linearity~of~expectation}$
@@
$=\left\langle\frac{\partial A(\boldsymbol{\phi})}{\partial\boldsymbol{\phi}},\boldsymbol{\phi}-\boldsymbol{\phi}^{\prime}\right\rangle+A(\boldsymbol{\phi}^{\prime})-A(\boldsymbol{\phi}),$ $\triangleright\text{Since~}\frac{\partial A(\boldsymbol{\phi})}{\partial\boldsymbol{\phi}}=\mathbb{E}_{\boldsymbol{x},\boldsymbol{z}\sim{p}_{\boldsymbol{\phi}}}\left[S(\boldsymbol{x},\boldsymbol{z})\right]$
@@
$\frac{1}{|\mathcal{D}|}\sum_{\boldsymbol{x}\in\mathcal{D}}\left\langle\frac{\partial\log{p}_{\boldsymbol{\phi}}(\boldsymbol{x})}{\partial\boldsymbol{\phi}},\boldsymbol{\phi}^{\prime}\right\rangle-\mathtt{KL}_{\boldsymbol{\phi}}(\boldsymbol{\phi}^{\prime}).$ (11)
@@
[…]
@@
$=\sum_{c^{\prime}\in\mathsf{ch}(n)}\exp(\phi_{n,c^{\prime}}),$ $\triangleright\text{Since~we~assume~}\forall c,Z_{c}(\boldsymbol{\phi})=1$
@@
$=\mathtt{TD}_{m}(n)\cdot\exp(\phi_{n,c})+\sum_{c^{\prime}\in\mathsf{ch}(n),c^{\prime}\neq c}\exp(\phi_{n,c^{\prime}}),$ $\triangleright\text{Since~}\forall c,\mathtt{TD}_{c}(c)=1$
@@
$=\sum_{b\in\mathcal{A}}\exp(\phi_{m,b})\cdot\mathtt{TD}_{b}(n).$ $\triangleright\text{Since~}\mathtt{TD}_{m}(b)=\exp(\phi_{m,b})$
@@
We move on to the second equality in Lemma 1. According to Vergari et al., (2021), $\mathtt{KL}_{\boldsymbol{\phi}}(\boldsymbol{\phi}^{\prime})$ can be computed recursively as follows (define $\mathtt{KL}_{\boldsymbol{\phi}}^{n}(\boldsymbol{\phi}^{\prime})$ as the KLD w.r.t. $n$):
[…]
@@
[…]
@@ B.3 Decomposition of Parameter Flows
$=\frac{\theta_{n,c}}{\tilde{{p}}_{\boldsymbol{\phi}}^{n}(\boldsymbol{x})}\cdot\frac{\partial\tilde{{p}}_{\boldsymbol{\phi}}^{n}(\boldsymbol{x})}{\partial\theta_{n,c}},$ $\triangleright\text{By~definition~}\theta_{n,c}=\exp(\phi_{n,c})$
@@
B.4 Computing TD-Prob Using Auto-Differentiation
In this section, we prove that $\mathtt{TD}(n)$ (and thus also $\mathtt{TD}(\phi_{n,c})$) can be computed by differentiating $Z_{n_{r}}(\boldsymbol{\phi})$, where $n_{r}$ is the root node. Note that the definition of $Z_{n}(\boldsymbol{\phi})$ follows Section B.1.
We proceed with the proof by induction. First, as a base case, we have that
$\frac{\partial Z_{n_{r}}(\boldsymbol{\phi})}{\partial Z_{n_{r}}(\boldsymbol{\phi})}=1=\mathtt{TD}(n_{r}).$
Next, assume that for a sum/input node $n$, for all its parents $m\in\mathsf{pa}(n)$ (which are product nodes according to Def. 1) we satisfy that
$\mathtt{TD}(m)=\frac{\partial Z_{n_{r}}(\boldsymbol{\phi})}{\partial Z_{m}(\boldsymbol{\phi})}.$
We proceed to derive $\partial Z_{n_{r}}(\boldsymbol{\phi})/\partial Z_{n}(\boldsymbol{\phi})$:
$\frac{\partial Z_{n_{r}}(\boldsymbol{\phi})}{\partial Z_{n}(\boldsymbol{\phi})}$ $=\sum_{m\in\mathsf{pa}(n)}\frac{\partial Z_{n_{r}}(\boldsymbol{\phi})}{\partial Z_{m}(\boldsymbol{\phi})}\cdot\frac{\partial Z_{m}(\boldsymbol{\phi})}{\partial Z_{n}(\boldsymbol{\phi})},$
$=\sum_{m\in\mathsf{pa}(n)}\mathtt{TD}(m)\cdot\frac{\partial Z_{m}(\boldsymbol{\phi})}{\partial Z_{n}(\boldsymbol{\phi})},$
$=\sum_{m\in\mathsf{pa}(n)}\mathtt{TD}(m).$ $\triangleright\text{By~definition~}\frac{\partial Z_{m}(\boldsymbol{\phi})}{\partial Z_{n}(\boldsymbol{\phi})}=1$
The final case is for a product node $n$, assuming all its parents satisfy the requirement, i.e., $\forall m\in\mathsf{pa}(n),\mathtt{TD}(m)=\partial Z_{n_{r}}(\boldsymbol{\phi})/\partial Z_{m}(\boldsymbol{\phi})$. We can simplify the gradient with respect to $Z_{n}(\boldsymbol{\phi})$ by
$\frac{\partial Z_{n_{r}}(\boldsymbol{\phi})}{\partial Z_{n}(\boldsymbol{\phi})}$ $=\sum_{m\in\mathsf{pa}(n)}\frac{\partial Z_{n_{r}}(\boldsymbol{\phi})}{\partial Z_{m}(\boldsymbol{\phi})}\cdot\frac{\partial Z_{m}(\boldsymbol{\phi})}{\partial Z_{n}(\boldsymbol{\phi})},$
$=\sum_{m\in\mathsf{pa}(n)}\mathtt{TD}(m)\cdot\frac{\partial Z_{m}(\boldsymbol{\phi})}{\partial Z_{n}(\boldsymbol{\phi})},$
$=\sum_{m\in\mathsf{pa}(n)}\mathtt{TD}(m)\cdot\theta_{m,n}.$ $\triangleright\text{By~definition~}\frac{\partial Z_{m}(\boldsymbol{\phi})}{\partial Z_{n}(\boldsymbol{\phi})}=\theta_{m,n}$
@@ Appendix C Global Parameter Renormalization of PCs
$\forall\boldsymbol{x},\;\tilde{{p}}_{\boldsymbol{\phi}^{\prime}}^{n}(\boldsymbol{x})=\frac{1}{Z_{n}(\boldsymbol{\phi})}\cdot\tilde{{p}}_{\boldsymbol{\phi}}^{n}(\boldsymbol{x}),$
where $Z(\boldsymbol{\phi}):=\sum_{\boldsymbol{x}}\tilde{{p}}_{\boldsymbol{\phi}}^{n}(\boldsymbol{x})$ is the partition function of $\tilde{{p}}_{\boldsymbol{\phi}}^{n}$.
@@
We begin by proving the correctness of the algorithm. Specifically, we show by induction that $\tilde{{p}}_{\boldsymbol{\phi}^{\prime}}^{n}(\boldsymbol{x})=\tilde{{p}}_{\boldsymbol{\phi}}^{n}(\boldsymbol{x})/Z_{\boldsymbol{\phi}}(n)$ for every $n$ and $\boldsymbol{x}$. In the base case, all input nodes satisfy the equation since they are assumed to be normalized. Next, given a product node $n$, assume the distributions encoded by all its children $c$ satisfy
$\forall c\in\mathsf{ch}(n),\;\tilde{{p}}_{\boldsymbol{\phi}^{\prime}}^{c}(\boldsymbol{x})=\tilde{{p}}_{\boldsymbol{\phi}}^{c}(\boldsymbol{x})/Z_{\boldsymbol{\phi}}(c).$ (18)
@@
$\tilde{{p}}_{\boldsymbol{\phi}^{\prime}}^{n}(\boldsymbol{x})$ $=\prod_{c\in\mathsf{ch}(n)}\tilde{{p}}_{\boldsymbol{\phi}^{\prime}}^{c}(\boldsymbol{x})=\prod_{c\in\mathsf{ch}(n)}\tilde{{p}}_{\boldsymbol{\phi}}^{c}(\boldsymbol{x})/Z_{\boldsymbol{\phi}}(c),$
$=\frac{\prod_{c\in\mathsf{ch}(n)}\tilde{{p}}_{\boldsymbol{\phi}}^{c}(\boldsymbol{x})}{\prod_{c\in\mathsf{ch}(n)}Z_{\boldsymbol{\phi}}(c)},$
$=\frac{\tilde{{p}}_{\boldsymbol{\phi}}^{n}(\boldsymbol{x})}{Z_{\boldsymbol{\phi}}(n)}.$
Finally, consider a sum node $n$ whose children satisfy Equation 18. We simplify $\tilde{{p}}_{\boldsymbol{\phi}^{\prime}}^{n}(\boldsymbol{x})$ in the following:
$\tilde{{p}}_{\boldsymbol{\phi}^{\prime}}^{n}(\boldsymbol{x})$ $=\sum_{c\in\mathsf{ch}(n)}\theta^{\prime}_{n,c}\cdot\tilde{{p}}_{\boldsymbol{\phi}^{\prime}}^{c}(\boldsymbol{x}),$
@@
$=\sum_{c\in\mathsf{ch}(n)}\frac{\theta_{n,c}\cdot\cancel{Z_{\boldsymbol{\phi}}(c)}}{Z_{\boldsymbol{\phi}}(n)}\cdot\frac{\tilde{{p}}_{\boldsymbol{\phi}}^{c}(\boldsymbol{x})}{\cancel{Z_{\boldsymbol{\phi}}(c)}},$ $\triangleright\text{By~induction}$
$=\frac{\sum_{c\in\mathsf{ch}(n)}\theta_{n,c}\cdot\tilde{{p}}_{\boldsymbol{\phi}}^{c}(\boldsymbol{x})}{Z_{\boldsymbol{\phi}}(n)},$
$=\tilde{{p}}_{\boldsymbol{\phi}}^{n}(\boldsymbol{x})/Z_{\boldsymbol{\phi}}(n).$
@@
$\forall\boldsymbol{x},\;\frac{\partial\log\tilde{{p}}_{\boldsymbol{\phi}^{\prime}}(\boldsymbol{x})}{\partial\phi^{\prime}_{n,c}}=\frac{\partial\log\tilde{{p}}_{\boldsymbol{\phi}}(\boldsymbol{x})}{\partial\phi_{n,c}}.$
@@
$\forall\boldsymbol{x},\;\frac{\partial\log\tilde{{p}}_{\boldsymbol{\phi}^{\prime}}^{n}(\boldsymbol{x})}{\partial\log\tilde{{p}}_{\boldsymbol{\phi}^{\prime}}^{c}(\boldsymbol{x})}=\frac{\partial\log\tilde{{p}}_{\boldsymbol{\phi}}^{n}(\boldsymbol{x})}{\partial\log\tilde{{p}}_{\boldsymbol{\phi}}^{c}(\boldsymbol{x})}.$
@@
$\log\tilde{{p}}_{\boldsymbol{\phi}^{\prime}}^{n}(\boldsymbol{x})=\sum_{c\in\mathsf{ch}(n)}\log\tilde{{p}}_{\boldsymbol{\phi}^{\prime}}^{c}(\boldsymbol{x}).$
Hence, its derivative w.r.t. $\log\tilde{{p}}_{\boldsymbol{\phi}^{\prime}}^{c}(\boldsymbol{x})$ is $1$.